OpenStudy (anonymous):

Calculate the requested derivative from the given information. Given f(u)=u^3and g(x) =u=x+4/x-2,find (f of g)'(3).

OpenStudy (anonymous):

knowing f(x) = x^3 and g(x) = x + 4/(x - 2), can you determine f(g(x)) ? if so, you can determine f(g(x))'

OpenStudy (anonymous):

Ok let me see

OpenStudy (anonymous):

So it would start off ... (x+4/x-2)^3

OpenStudy (anonymous):

right. is that \[x + 4/x - 2\] or \[x + \frac{ 4 }{ x - 2 }\] ?

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

ah ok.

OpenStudy (anonymous):

So next would be ....... 3(x+4/x−2)

OpenStudy (anonymous):

yes you have to determine the following derivative: \[\frac{ d }{ dx } (x + \frac{ 4 }{ x } - 2)^3\]

OpenStudy (anonymous):

Huh? I don't know how you got that

OpenStudy (anonymous):

do you agree that\[f(g(x)) =( x + \frac{ 4 }{ x } - 2)^3\] ?

OpenStudy (anonymous):

No, how did it come from x+4 to (x+4x−2)3 _____ x-2

OpenStudy (anonymous):

you stated\[g(x) = x + \frac{ 4 }{ x } - 2\] ?

OpenStudy (anonymous):

no, g(x) is x+4 x-2

OpenStudy (anonymous):

\[f(g(x)) = \left( \frac{ x + 4 }{ x - 2 } \right)^3\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah ok. so we just need to take the derivative of that thing. then evaluate it at x = 3

OpenStudy (anonymous):

3 * ( (x+4/x−2)^ 1/2 ?

OpenStudy (anonymous):

i get \[3\left( \frac{ x + 4 }{ x - 2 } \right)^2 * \frac{ (x - 2) - (x + 4) }{ (x - 2)^2 }\]

OpenStudy (anonymous):

using chain rule, then quotient rule

OpenStudy (anonymous):

ok, let me see, what I get

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

so how is the answer -882

OpenStudy (anonymous):

i got the same answer

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

let x = 3, and solve

OpenStudy (anonymous):

\[f(g(x))' = 3\left( \frac{ x+4 }{ x - 2 } \right)^2 * \frac{ (x - 2) - (x + 4)}{ (x - 2)^2 }\] \[f(g(3))' = 3\left( \frac{ 3+4 }{ 3 - 2 } \right)^2 * \frac{ (3 - 2) - (3 + 4) }{ (3 - 2)^2 }\] \[ = 3*7^2 * -6 = -882\]

OpenStudy (anonymous):

oh, I didn't square the 7 at first. but i got it now , thans